OpenStudy (anonymous):

What is the value of the expression 4/2^2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oh my bad the question is masted up

OpenStudy (anonymous):

(4/2)^2 is really just 2^2, which equals 4.

OpenStudy (anonymous):

it is suppost to be 4/2^-2

OpenStudy (anonymous):

Then the answer is 1/4 instead.

OpenStudy (anonymous):

(4/2)^-2 = (2^-2) = (1/2^2) = (1/4)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Whenever you see a negtive exponant, it means the answer is going to be 1/something

OpenStudy (anonymous):

like 2^-2 = 1/4

OpenStudy (anonymous):

I got 8

OpenStudy (anonymous):

can you type out your work?

OpenStudy (anonymous):

ok 4/2^-2 = 4 * 1/2^-2 = 4 * 2^2 = 4 * 4 =8

OpenStudy (anonymous):

Ok. First thing is to remember the order of operations. (4/2)^-2 should be in parenthesis, so you reduce inside the parenthesis first.

OpenStudy (anonymous):

So (4/2)^-2 = 2^-2

OpenStudy (anonymous):

there are no parenthesis

OpenStudy (anonymous):

They are understood to be there even if not written.

OpenStudy (anonymous):

just like the expression x+2/x-2 may not show parenthesis

OpenStudy (anonymous):

but it is understood to be (x+2)/(x-2)

OpenStudy (anonymous):

ok but would u say the way i did was correct

OpenStudy (anonymous):

just one of those math things you have to look out for

OpenStudy (anonymous):

No

OpenStudy (anonymous):

If you dont follow the proper order of operations you wont get the correct answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you remember the order of operations?